OpenStudy (anonymous):

Please help! The perimeter of a triangle is 39 inches. If the length of the shortest side is 1/2 the length of the longest side, and the length of the third side is 1 less the length of the longest side, what is the length of each side? If x is the shortest side and y is the longest side, one equation is...?

OpenStudy (anonymous):

Here are all of the equations you can get. Let us set x as the short side, y as the longest side, and z as the third side. 1. x + y + z = 39 2. x = 1/2 y 3. z = y - 1

OpenStudy (anonymous):

From these three equations, you should be able to find the equation.

OpenStudy (anonymous):

ok so take the longest side to be \[x.\] The shortest side will be \[1/2(x)\] The 3rd side will be\[ x-1\] \[So x+1/2x+(x-1)=39 inches\]

OpenStudy (anonymous):

Thanks guys! Wish I could give more than one medal :(

OpenStudy (anonymous):

No problem.